OpenStudy (anonymous):

A bicyclist travels at a constant rate on her bicycle. Biking with the wind, it takes 2.6 hours to travel 52 miles. Biking against the wind, the return trip takes 4 hours. If there is no wind, how fast was the cyclist biking? (step by step help, I tried it multiple times but got lost and confused completely)

OpenStudy (anonymous):

you are looking for the rate yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

one way is to use distance equals rate time time the rate going is \(r+w\) and the rate returning is \(r-w\) then \[52=(r+w)\times 2.6\] and \[52=(r-w)\times4\]

OpenStudy (anonymous):

where \(r\) is the rate you are looking for and \(w\) is the rate of the wind

OpenStudy (anonymous):

ok but then how do I solve it from there?